# Value function interference and greedy action selection in value-based multi-objective reinforcement learning

### 09 Feb 2024 v1

@@ 3 $a_{2}$ $a_{2}$ (0, -5, -5) -25
This example illustrates the core concept of value function interference in value-based multi-objective RL. If the utility function allows actions with widely differing vector outcomes to map to the same utility value, then the vector value function learned for earlier states may be inconsistent with the actual optimal policy. We note that if the utility function is linear then value interference does not impact on selecting the optimal action. The value learned for $Q(A,a_{1})$ will lie (approximately) on the line segment joining between $Q(B,a_{1})$ and $Q(B,a_{2})$. If $Q(B,a_{1})$ and $Q(B,a_{2})$ share the same utility value for a particular linear weighting of objectives, then all points on that line also have that same utility.
2.2 Empirical evaluation of a potential solution
The problem of value function interference in a stochastic MOMDP can potentially be addressed by modifying the selection process so that ties between multiple greedy actions are broken in a deterministic fashion. For example, we might always select the greedy action with the lowest sub-script - this introduces some degree of bias in the action selection, but presumably the user is indifferent to this bias (otherwise it would be represented in some way within the utility function).
To demonstrate the impact of value function interference, and the ability of non-random tie-breaking to mitigate its effects, empirical experiments were run on the MDP from Figure 1. These experiments used three variations of a tabular multi-objective Q-learning agent. The first variant is a baseline using the standard random approach to tie-breaking in the case of actions with equal scalarised values. The second variant deterministically breaks ties by favouring the action with the lower index (i.e. $a_{1}$ is preferred to $a_{2}$). As this could produce a bias towards the desired policies (both of which select $a_{1}$ in the first state), we also included a third variant which favours the action with the higher index.
Each agent’s performance was evaluated over a hyperparameter sweep across different values for the learning rate ($\alpha$) and starting exploration parameter (using e-greedy exploration with $\epsilon$ linearly decayed to zero over the trial). The $\lambda$ and $\gamma$ hyperparameters were fixed at 0.95 and 1 respectively, and Q-values were optimistically initialised to (12,0,0). For each set of hyperparameters 100 trials of each agent were executed. Each trial lasted for 500 episodes, and at the end of the trial the final greedy policy of the agent was identified.
@@ 2.2.1 Random tie-breaking
The results from the random tie-breaking agent (Figure 2, top row) clearly illustrate the problems caused by value function interference, as for large areas of the hyperparameter sweep the agents converged to the sub-optimal Policy 2 in the overwhelming majority of trials. Notably this was most likely to occur for trials where the learning rate $\alpha$ was in the middle of its range $0..1$.
This can be explained by the nature of the utility function. The value learned for $Q(A,a_{1})$ will over time converge to lie on the line segment between (7,-1,-5) and (7, -5, -1) (i.e. between the rewards for $B,a_{1}$ and $B,a_{2}$. The location along this line will depend on the frequency with which each action has been performed in state $B$ recently, as well as the magnitude of $\alpha$. Points on this line-segment will have the form $p=(7,-5+4x,-1-4x)$ where $0\leq x\leq 1$. Action $a_{1}$ will be preferred to $a_{2}$ in state $A$ when $U(Q(A,a_{1}))-U(Q(A,a_{2}))>0$. Substituting the definition of $p$ and the value of $Q(A,a_{2})$ yields that this will occur when $016x^{2}-16x+2>$ 0, which occurs when $x<(2-\sqrt{(}2))/4$ and when $x>(2+\sqrt{(}2))/4$ (i.e. when $Q(A,a_{1})$ is close to either $Q(B,a_{1})$ or $Q(B,a_{2})$.
Consider the situation when $Q(A,a_{1})$ is a good approximation to $Q(B,a_{1})$, and now action $B,a_{2}$ is selected as a greedy action and executed. $Q(A,a_{1})$ will be updated towards (7, -5, -1). A small value of $\alpha$ will produce a small change in $Q(A,a_{1})$ meaning it will likely still be in the range where it is considered preferable to $Q(A,a_{2})$. A large value of $\alpha$ will produce a large change in $Q(A,a_{1})$, jumping over the portion of the line-segment where $A,a_{2}$ is preferred. Meanwhile middling values of $\alpha$ will result in $Q(A,a_{1})$ falling roughly halfway between (7,-1,-5) and (7, -5, -1) (i.e. $x\approx 0.5$), which is the region where $A,a_{2}$ will be preferred. Hence it is this central range of $\alpha$ values where convergence to the incorrect policy is most likely to occur, which is evident in the experimental results.
There is also a tendency for sub-optimal convergence to become more frequent across a wider range of $\alpha$ values as the degree of exploration is increased. We note that the exploration values used here may seem high for such a simple MOMDP, but past research has shown that exploration rates may need to be higher for MORL than for single-objective RL (Vamplew et al., 2017b). Therefore this link between value function interference and high rates of exploration may prove to be problematic for more complex MOMDPs.
@@ 2.2.2 Deterministic tie-breaking
Given this capacity for the deterministic tie-breaking to induce different preferences over policies, we need to consider the possibility that the reduction in the frequency of convergence to sub-optimal policies is not due directly to the elimination of random tie-breaking, but instead due to an explicit bias away from Policy 2. Figure 3 assists in differentiating between these causes by performing a pairwise comparison of the agents based on how frequently they incorrectly converge to Policy 2. It can be seen from Figure 3(left) that the low-index deterministic tie-breaking agent converges to this policy with lower or comparable frequency to the random tie-breaking agent for every hyperparameter setting which was tested. Overall it selects this policy in 1493 fewer trials than the random tie-breaking agent (almost 30% of the total number of trials). The high-index deterministic tie-breaking agent also improves substantially over the random tie-breaking agent, selecting the sub-optimal policy in 1385 fewer cases, but this is 108 cases worse than the low-index agent (Figure 3 (middle and right)). So while the low-index agent’s bias away from action $a_{2}$ in state $A$ is partially responsible for the reduction in trials converging to Policy 2, this accounts for just 108/1493 cases – more than 90% of the improvement is due to avoiding the stochasticity of the random tie-breaker rather than the specific bias of the deterministic tie-breaker.
@@
Figure 3: Heatmaps showing the difference in the frequency with which the incorrect Policy 2 was selected as the final greedy policy across 100 trials using different hyperparameter settings for each combination of the three different tie-breaking approaches: Higher-index -random (top row); Lower-indexed - random (middle row); lower-index - higher-indexed (bottom row)
@@
These results demonstrate that deterministic tie-breaking may be beneficial in the case of MOMDPs where some actions share identical utility. However more generally the actions need not map to exactly the same scalar utility value in order for interference to arise. In practice the Q-value estimates may be noisy due to either the use of function approximation and/or stochasticity in the environment. These errors in estimates may result in changes in the action which is selected as greedy in each state, leading to variations over time in the vector value updates for prior states, which could result in value function interference. This is particularly likely to occur for highly non-linear definitions of utility as small variations in the estimated vector values can lead to large variations in the estimated utility; for example, we have previously observed unstable and slow learning due to value function interference when using a thresholded lexicographic approach to utility on stochastic problems (Vamplew et al., 2020, 2021; Ding et al., 2024).
3 Interference in multi-objective value functions for stochastic environments
If we now consider environments with some element of stochasticity (in either the starting states, state transitions or rewards), then the SER and ESR criteria are no longer equivalent, and value interference issues may arise for ESR-optimal learning which do not occur for SER-optimal learning.
Consider the very simple MOMDP  [^3] In fact, as this MOMDP has only a single non-terminal state, this example demonstrates that value function interference can arise even for ESR optimisation of multi-objective multi-armed bandits (MOMABs). shown in Figure 4. Action $a_{1}$ has a stochastic reward which varies between (7, -1, -5) and (7, -5, -1) with equal probability. The utility of this action differs depending on whether we are optimising with respect to ESR or SER – assuming the same definition of utility as in Section 2, $U_{SER}(a_{1})$=5 (the scalarised value of the expected vector reward) while $U_{ESR}(a_{1})$=9 (the expected scalarised value, which is 9 for both of the possible outcomes). Action $a_{2}$ deterministically returns a reward of (8, -3, -3) – as this is deterministic $U_{SER}(a_{2})=U_{ESR}(a_{2})$ = 7.
Therefore the SER-optimal policy in this case is to select $a_{2}$, while the ESR-optimal policy selects $a_{1}$. The SER-optimal policy can be determined on the basis of the vector Q-values learned by the agent, but this is not true for the ESR-optimal policy, as these values only represent the expected vector return which is insufficient for the agent to determine the per-episode expected scalarised return. The use of deterministic tie-breaking proposed as a mitigating solution earlier in Section 2 is not useful in this context, as the stochasticity in the returns arises from the environment rather than being induced by stochasticity in the policy due to tie-breaking.
For this scenario we suggest that the most promising course of action is to incorporate concepts from distributional RL into MORL, as pioneered by Hayes et al. (2021); Reymond et al. (2023); Röpke et al. (2023). If the agent learns a distribution over the vector returns for each state-action, then this can be combined with the utility function so as to estimate the ESR. For example if the agent has correctly learned that the distribution of future returns for action $a_{1}$ is (-7,-1, -5) with $p=0.5$ and (-7,-5, -1) with $p=0.5$. From this information the agent can correctly estimate the ESR utility of $a_{1}$ as the probabilistically-weighted sum of those return vectors.
$\begin{split}U_{ESR}(a_{1})=0.5U_{ESR}(7,-1,-5)+0.5U_{ESR}(7,-5,-1)\\ =0.5*9+0.5*9\\ =9\end{split}$ (3)

### 22 Apr 2026 v2

@@ 3 $a_{2}$ - $a_{2}$ (0, -5, -5) -25
This example illustrates one of the ways that value function interference may arise within a deterministic environment. If the utility function allows actions with widely differing vector outcomes to map to the same utility value, then the vector value function learned for earlier states may be inconsistent with the actual optimal policy. We note that if the utility function is linear then value interference does not impact on selecting the optimal action. The value learned for $Q(A,a_{1})$ will converge towards a point on the line segment joining $Q(B,a_{1})$ and $Q(B,a_{2})$. If $Q(B,a_{1})$ and $Q(B,a_{2})$ share the same utility value for a particular linear weighting of objectives, then all points on that line also have that same utility.
2.3 Empirical evaluation of the effect of value-function interference in a deterministic MOMDP
The problem of value function interference in a deterministic MOMDP can potentially be addressed by modifying the selection process so that ties between multiple greedy actions are broken in a deterministic fashion. For example, we might always select the greedy action with the lowest sub-script - this introduces some degree of bias in the action selection, but presumably the user is indifferent to this bias (otherwise it would be represented in some way within the utility function).
To demonstrate the impact of value function interference, and the ability of non-random tie-breaking to mitigate its effects, empirical experiments were run on the MOMDP from Figure 4. These experiments used three variations of a tabular multi-objective Q-learning agent. The first variant is a baseline using the standard random approach to tie-breaking in the case of actions with equal scalarised values. The second variant deterministically breaks ties by favouring the action with the lower index (i.e. $a_{1}$ is preferred to $a_{2}$). As this could produce a bias towards the desired policies (both of which select $a_{1}$ in the first state), we also included a third variant which favours the action with the higher index.
Each agent’s performance was evaluated over a hyperparameter sweep across different values for the learning rate ($\alpha$) and starting exploration parameter (using $\epsilon$-greedy exploration with $\epsilon$ linearly decayed to zero over the trial). The $\lambda$ and $\gamma$ hyperparameters were fixed at 0.95 and 1 respectively, and to produce more directed exploration Q-values were optimistically initialised to (12,0,0). For each set of hyperparameters 100 trials of each agent were executed. Each trial lasted for 500 episodes, and at the end of the trial the final greedy policy of the agent was identified.
@@ 2.3.1 Random tie-breaking
The results from the random tie-breaking agent (Figure 5, top row) clearly illustrate the problems caused by value function interference, as for large areas of the hyperparameter sweep the agents converged to the sub-optimal Policy 2 in the overwhelming majority of trials. Notably this was most likely to occur for trials where the learning rate $\alpha$ was in the middle of its range $0\ldots 1$.
This can be explained by the nature of the utility function. The value learned for $Q(A,a_{1})$ will over time converge to lie on the line segment between (7,-1,-5) and (7,-5,-1) (i.e. between the rewards for $B,a_{1}$ and $B,a_{2}$. The location along this line will depend on the frequency with which each action has been performed in state $B$ recently, as well as the magnitude of $\alpha$. Points on this line-segment will have the form p=(7,5+4x,14x)p=(7,-5+4x,-1-4x) where $0\leq x\leq 1$. Action $a_{1}$ will be preferred to $a_{2}$ in state $A$ when $U(Q(A,a_{1}))-U(Q(A,a_{2}))>0$. Substituting the definition of $p$ and the value of $Q(A,a_{2})$ yields that this will occur when $16x^{2}-16x+2>0$, which occurs when $x<(2-\sqrt{(}2))/4$ and when $x>(2+\sqrt{(}2))/4$ (i.e. when $Q(A,a_{1})$ is close to either $Q(B,a_{1})$ or $Q(B,a_{2})$).
Consider the situation where $Q(A,a_{1})$ is a good approximation to $Q(B,a_{1})$, and now action $B,a_{2}$ is selected as a greedy action and executed. $Q(A,a_{1})$ will be updated towards (7,-5,-1). A small value of $\alpha$ will produce a small change in $Q(A,a_{1})$ meaning it will likely still be in the range where it is considered preferable to $Q(A,a_{2})$. A large value of $\alpha$ will produce a large change in $Q(A,a_{1})$, jumping over the portion of the line-segment where $A,a_{2}$ is preferred. Meanwhile middling values of $\alpha$ will result in $Q(A,a_{1})$ falling roughly halfway between (7,-1,-5) and (7, -5, -1) (i.e. $x\approx 0.5$), which is the region where $A,a_{2}$ will be preferred. Hence it is this central range of $\alpha$ values where convergence to the incorrect policy is most likely to occur, which is evident in the experimental results.
There is also a tendency for sub-optimal convergence to become more frequent across a wider range of $\alpha$ values as the degree of exploration is increased. We note that the exploration values used here may seem high for such a simple MOMDP, but past research has shown that exploration rates may need to be higher for MORL than for single-objective RL (Vamplew et al., 2017a). Therefore this link between value function interference and high rates of exploration may prove to be problematic for more complex MOMDPs.
@@ 2.3.2 Deterministic tie-breaking
Given this capacity for the deterministic tie-breaking to induce different preferences over policies, we need to consider the possibility that the reduction in the frequency of convergence to sub-optimal policies is not due directly to the elimination of random tie-breaking, but instead due to an explicit bias away from Policy 2. In particular low-index tie-breaking will tend to favour $a_{1}$ in State $A$, which may in itself reduce the likelihood of convergence to Policy 2, regardless of any impact on value function interference. Figure 6 assists in differentiating between these causes by performing a pairwise comparison of the agents based on how frequently they incorrectly converge to Policy 2. It can be seen from Figure 6(middle) that the low-index deterministic tie-breaking agent converges to this policy with lower or comparable frequency to the random tie-breaking agent for every hyperparameter setting which was tested. Overall it selects this policy in 1493 fewer trials than the random tie-breaking agent (almost 30% of the total number of trials). The high-index deterministic tie-breaking agent also improves substantially over the random tie-breaking agent, selecting the sub-optimal policy in 1385 fewer cases (Figure 6 (left), but this is 108 cases worse than the low-index agent. So while the low-index agent’s bias away from action $a_{2}$ in state $A$ is partially responsible for the reduction in trials converging to Policy 2, this accounts for just 108/1493 cases – more than 90% of the improvement is due to avoiding the stochasticity of the random tie-breaker rather than the specific bias of the deterministic tie-breaker.
@@
Figure 6: Heatmaps showing the difference in the frequency with which the incorrect Policy 2 was selected as the final greedy policy across 100 trials using different hyperparameter settings for each combination of the three different tie-breaking approaches: Higher-index - Random (left); Lower-index - Random (middle); Lower-index - Higher-index (right)
These results demonstrate that deterministic tie-breaking may be beneficial in the case of MOMDPs where some actions share identical utility.
2.4 Interference due to changes in greedy action
The results in the previous section demonstrate that non-random tie-breaking can alleviate the impact of value function interference. However it fails to eradicate this problem, as there are still a considerable number of cases where the agent converges to the non-optimal Policy 2 (including some parameter settings where this occurs in all trials). Therefore there must be a secondary cause of value function interference within this deterministic MOMDP.
We hypothesise that this may occur in cases where the agent’s choice of greedy action changes. In a deterministic MOMDP this could arise due to optimistic initialisation of the Q-values – each action will need to be sampled multiple times before its associated Q-value estimates converge closely enough to their true values for the optimality (or otherwise) of that action to be accurately established. The number of samples required will depend on the learning rate, and also on the size of the gap in utility between each action and the optimal action.
Consider a variant of the MOMDP from Figure 4, in which the reward vector for $a_{2}$ in state B is $(7-\delta,-5,-1)$ meaning that its utility is now lower than that of action $a_{1}$ by $2\delta$. This change means that Policy 0 ($a_{1}$$a_{1}$) is now the sole optimal policy. The magnitude of $\delta$ will inversely influence the number of samples and Q-value updates required to establish that $U(Q(B,a_{1}))>U(Q(B,a_{2}))$, and hence will reduce the number of occasions on which the agent switches its choice of greedy action for state $B$. Note that this change to the environment has no impact on the relative utility of Policy 0 ($a_{1}$$a_{1}$) and Policy 2 ($a_{2}$$a_{1}$) and so, absent the effects of value function interference, changing the value of $\delta$ should have no effect on how frequently Policy 2 is selected by the agent.
Original MOMDP
$\delta$=0.0001
$\delta$=0.1
Figure 7: Heatmaps showing the frequency with which the correct Policy 0 and incorrect Policies 1 and 2 were selected as the final greedy policy for the deterministic MOMDP across 100 trials using different hyperparameter settings. The top row corresponds to the original MOMDP from Figure 4, while the other rows use a modified reward for action $a_{2}$ in state $B$, where the first objective has been reduced by $\delta$. These experiments use the random tie-breaker.
However the results in Figure 7 show that even a tiny $\delta$ value of 0.0001 is sufficient to allow the agent to converge to the optimal policy (Policy 0) in almost all trials across all values of the learning hyperparameters. A larger $\delta$ value is sufficient for all trials to converge successfully to this optimal policy.
Original
$\delta$=0.0001
$\delta$=0.1
Figure 8: Heatmaps showing the frequency with which each of the policies were selected during learning, as a percentage of all episodes over all 100 trials. The top row corresponds to the original MOMDP from Figure 4, while the other rows use a modified reward for action $a_{2}$ in state $B$, where the first objective has been reduced by $\delta$. These experiments use the random tie-breaker.
Further insight into the effect of $\delta$ on the prevalence of value function interference can be gained by looking at the selection of actions during learning rather than just in the final policy. Figure 8 shows the percentage of episodes in which each policy is executed during learning (including episodes in which exploratory actions are selected). It can be seen that as the magnitude of $\delta$ increases, the agent increasingly favours the optimal Policy 0 during training. Significantly, this is due not just to less frequent execution of Policy 1 (whose utility is directly effected by $\delta$) but also by reduced execution of Policies 2 and 3 – this provides clear evidence that value function interference must be arising in the estimated value of $Q(A,a_{1})$ for smaller values of $\delta$.
2.5 Value function interference – summary
The results reported in this section establish the existence of the learning phenomenon that we have named value function interference. This arises when the expected vector returns learned by an RL agent are inadequate to effectively determine the actions and policies which are optimal with respect to the utility function. This can significantly hinder learning, and may lead to convergence to a sub-optimal policy.
The most obvious situation in which value function interference can occur is where the environment itself is stochastic with regards to either rewards and/or state transitions. However we have demonstrated that interference can also occur in deterministic environments, either due to stochasticity in tie-breaking of greedy actions, or due to frequent changes in the choice of greedy action (i.e. as might be induced by optimistic initialisation). The latter issue may be especially problematic in the context of value-based deep RL. Schaul et al. (2022) identified extremely high rates of change in the greedy policy in value-based deep RL algorithms such as DQN and R2D2, which suggests that multi-objective extensions of those algorithms may be particularly susceptible to value-function interference.
@@ 2.5 Value function interference – summary
The experiments in this section demonstrated value function interference arising even in the context of otherwise stable learning algorithms based on tabular representation of Q-values. In practice for more complex environments the Q-value estimates may themselves be noisy due to the need to use function approximation. These errors in estimates may result in changes in the action which is identified as greedy in each state, leading to variations over time in the vector value updates for prior states, which could result in increased levels of value function interference. This is particularly likely to occur for highly non-linear definitions of utility as small variations in the estimated vector values can lead to large variations in the estimated utility; for example, unstable and slow learning has previously been reported when using a thresholded lexicographic approach to utility on stochastic problems (Vamplew et al., 2021a; b; Ding et al., 2025).
3 Sensitivity of MORL to overestimation
It is well established that off-policy methods such as Q-learning are prone to overestimation of Q-values when used in conjunction with function approximation, and there has been considerable effort at addressing this issue (Anschel et al., 2017; Van Hasselt et al., 2018; Ly et al., 2024). Li (2023) notes that for single-objective RL, overestimation may not actually prevent the optimal policy from being learned as “if all values are uniformly higher, relative action preferences are still preserved”. More formally:
$Q(s,a^{*})>Q(s,a)\implies Q(s,a^{*})+\delta>Q(s,a)+\delta$ (3)
While this would also apply to MORL with linear scalarisation, it is not necessarily true if non-linear scalarisation is used (Giuseppi and Pietrabissa, 2020; Mabsout et al., 2025). For a non-linear utility function, even a small amount of overestimation in the Q-value vectors may be sufficient to change the preference ordering of actions, even if that overestimation is consistent over all actions:
$U(Q(s,a^{*}))>U(Q(s,a))\mathrel{{\vtop{\halign{#\cr$\not\phantom{=}$\cr$\implies$\crcr}}}}U(Q(s,a^{*})+\delta)>U(Q(s,a)+\delta)$ (4)
In the remainder of this section, we present a brief empirical demonstration of the impact which overestimation, consistent or otherwise, might have on the behaviour of a value-based MORL agent.
3.1 Design of overestimation experiments
The benchmark environment for our experiments with overestimation is the DST-Mixed task from Vamplew et al. (2015) (Figure 9). This is a small grid-world where the agent controls the movements of a submarine searching for undersea treasure. There are ten Pareto-optimal deterministic policies representing different trade-offs between the time and treasure objectives. The Pareto front contains several concavities – as a result only four of these policies can be optimal under linear scalarisation.
Figure 9: A visualisation of the DST-Mixed environment (left), and its Pareto front (right). Black points show policies which can be found by linear scalarisation, while grey points are policies which can only be found via non-linear scalarisation. (Reproduced from Vamplew et al. (2015))
For simplicity we map this environment to a multi-objective multi-armed bandit (MOMAB), where each arm of the bandit corresponds to one of the Pareto-optimal policies. This allows us to avoid the additional complexities arising from variations in overestimation at different states within a trajectory. We also omit the process of learning Q-values. Instead we focus on the impact of overestimation of the Q-values for the starting state on the choice of policy to be executed, and how this varies depending on the nature of the utility function. Three utility functions were considered – linear scalarisation, thresholded lexicographic ordering (TLO), and Chebyshev distance. The latter uses a utopian reference point $z^{*}$. In the MORL literature, this is typically derived from the maximum return received for each objective during learning (i.e. this value is updated during the learning process) (Van Moffaert et al., 2013; Qin et al., 2020; Kim et al., 2022). As these experiments were not simulating the learning process itself, this point was set to the maximum value of each objective under any policy, plus a small positive offset so $z^{*}=(20.1,-0.9)$.
$U_{linear}(r_{treasure},r_{time})=w_{1}r_{treasure}+w_{2}r_{time}$ (5)
$U_{TLO}(r_{treasure},r_{time})=\begin{cases}r_{treasure},&\text{if }r_{time}\geq\text{threshold}\\ r_{time},&\text{otherwise}\end{cases}$ (6)
[…]
Utility function U Parameters Range
Linear Objective weights $w_{1}$, $w_{2}$ $w_{1}$ in $0.17\ldots 0.65{]}$, $w_{2}$ = 1 - $w_{1}$
TLO Threshold for time $20\ldots 0{]}$
Chebyshev distance Objective weights $w_{1}$, $w_{2}$ $w_{1}$ in $0.01\ldots 0.99{]}$, $w_{2}$ = 1 - $w_{1}$
Table 2: The ranges used for the utility function parameters in the overestimation experiments. Values were sampled uniformly within these ranges.
Prior to each arm-pull the following process was performed:
  • A random set of parameters were generated for the utility function. These were drawn uniformly from the ranges specified for each utility function in Table 2. These ranges were chosen as they give rise to an approximately even distribution over the policies when applied to the actual Q-values.
  • Positive noise was temporarily added to the true Q-values for each arm to simulate the effect of overestimation.
  • The arm to pull was selected greedily based on these ‘overestimated’ Q-values.
This process was repeated 1000 times for each utility function. The impact of the overestimation was tracked via the following statistics:
  • The number of pulls for which a non-optimal arm was selected.
  • The mean normalised regret incurred. For any arm pull, the regret can be calculated in terms of the loss in utility incurred by selecting this arm rather than the optimal arm
    $regret(a)=E[U(r_{a^{*}})]-E[U(r_{a})]\text{, where }a^{*}\text{ is the optimal arm}$ (8)
    However the range of the utility varies substantially between the three different utility functions, as can be seen in the colorbars in Figure 11, so a direct comparison of regret is not appropriate. Therefore we use a worst-case normalised regret measure, where the regret from the performed action is normalised with respect to the regret of the worst-possible action. This gives a value in the range $0\ldots 1$, enabling comparison across utility functions with different ranges.
    $regret_{WCN}(a)=regret(a)/regret(a^{worst})\text{, where }a^{worst}\text{ is the least-optimal arm}$ (9)
The entire experimental process was repeated twice for two different types of overestimation. In the consistent overestimation trials the same overestimation value $\delta$ was added on to both objectives for all Q-values. In the inconsistent overestimation trials for each arm a different overestimation value was uniformly sampled from the range $0\ldots\delta$ for each objective. In both cases multiple experiments were run for values of $\delta$ between 0 and 5.
3.2 Overestimation experiments results and discussion
Figure 10 illustrates the impact of consistent overestimation. As expected, this has no impact on the performance of an agent using linear scalarisation, regardless of the magnitude of the overestimation. The Q-values of all arms are increased by the same amount which preserves their ordering. In contrast the highly non-linear nature of TLO makes it very sensitive to overestimation. Even a minimal level of overestimation is sufficient to cause a substantial number of incorrect decisions. The number of errors increases rapidly as the magnitude of overestimation increases, until almost all decisions are incorrect. Meanwhile the other non-linear utility function Chebyshev distance performs somewhere between these extremes. Examining the normalised regret metric, it can be seen that not only is TLO prone to a higher frequency of errors than Chebyshev, but the impact of those errors (in terms of lost utility) is also much greater.
Figure 10: Results for the DST-Mixed environment for various utility functions with consistent overestimation of Q-values. The left plot shows the number of incorrect policy choices over 1000 trials, while the right plot shows the mean normalised loss of utility.
The utility contour plots in Figure 11 provide insight into the reason for these results. It can be seen that the gradient of the linear utility function is constant across the entire objective space. Therefore adding a constant level of overestimation to all Q-values increases the perceived utility of all policies equally and so has no impact on action selection. The Cheybshev function has a similarly constant gradient for large parts of the objective space, but the use of the max operator in conjunction with objective weights can result in inconsistent gradients. For example in the case shown in Figure 11, the optimal policy based on the true Q-values is the 7th from the left (#7), whereas once consistent noise is added the selected policy will be the 6th from the left (#6). Meanwhile the TLO utility function exhibits a uniform gradient if time exceeds the threshold, but a discontinuity at that threshold. Therefore adding constant overestimation results in varying changes in utility between the different policies – in particular the policy with true return (18.5, -14) undergoes a marked increase in utility when the overestimation is added, jumping from the 8th-ranked policy to the 1st-ranked. Where a policy’s return lies just below the threshold value, even a small amount of overestimation can result in incorrect action selection and a large loss in utility.
Figure 11: Contour plots mapping treasure/time to utility for a representative choice of each utility function: (top left) linear scalarisation with $w_{1}=0.65,w_{2}=0.35$, (top right) TLO with a time threshold of -13, and (bottom) Chebyshev distance with $w_{1}=0.65,w_{2}=0.35$. Black dots show the true return for the Pareto-front of policies. Arrows and red dots show the addition of consistent noise of +2.5 added to those true returns to represent over-estimated Q-values.
The results of the experiments using inconsistent overestimation are shown in Figure 12. For all three utility functions, inconsistent overestimation is more disruptive than consistent overestimation. [^5] When comparing Figures 10 and 12, keep in mind that for the former the level of overestimation is always equal to $\delta$, whereas for the latter the overestimation lies between 0 and $\delta$ and so will average $\delta/2$. So, for example, the most appropriate comparison for results at $\delta=0.2$ in Figure 10 are the results for $\delta=0.4$ in Figure 12. In this case, linear scalarisation is also affected as the variations in overestimation between different arms can cause the preferences between those arms to change. Nevertheless it still remains considerably more robust than TLO, both in terms of the rate of errors and the utility lost due to those errors. Interestingly while Chebyshev distance is less robust than linear scalarisation with regards to the number of incorrect policy selections, it does seem to be slightly more robust with regards to normalised regret for larger magnitudes of inconsistent overestimation.
Figure 12: Results for the DST-Mixed environment for various utility functions with inconsistent overestimation of Q-values. The left plot shows the number of incorrect policy choices over 1000 trials, while the right plot shows the mean normalised loss of utility.